A car traveling at 60 km/h will skid 30 m when its brakes are locked. If the same car is traveling at 180 km/h, what will be the skidding distance? Be sure to show all work to support your answer.
Answer by addingup(3677)

The distance is proportional to the square of the speed. Since 180/60 = 3 and 3^2 = 9, the distance will be 30*9
:
NOTE: Based on Kinetic Energy - Coefficient of Kinetic Friction:
friction*distance = 1/2mv^2 so distance is proportional to v^2